company can produce and sell x-units of one commodity and y-units of other commodity

at a pro€t given as P(x, y) = 400x + 500y − x

2 − y

2 − xy − 20000. Find the units and amount

for which the pro€t is maximum and minimum.



1
Expert's answer
2021-10-25T16:39:25-0400
"P(x, y)=400x+500y-x^2-y^2-xy-20000"

"\dfrac{\partial P}{\partial x}=400-2x-y"

"\dfrac{\partial P}{\partial y}=500-2y-x"

Find the critical point(s)


"\begin{cases}
 \dfrac{\partial P}{\partial x}=0 \\
\\
 \dfrac{\partial P}{\partial y}=0 
\end{cases}=> \begin{cases}
 400-2x-y=0 \\
\\
 500-2y-x=0 
\end{cases}"

"=> \begin{cases}
 y=400-2x \\
\\
 500-800+4x-x=0 
\end{cases}=> \begin{cases}
 x=100 \\
\\
 y=200 
\end{cases}"



"\dfrac{\partial^2 P}{\partial x^2}=-2"

"\dfrac{\partial^2 P}{\partial x \partial y}=\dfrac{\partial^2 P}{\partial y \partial x}=-1"


"\dfrac{\partial^2 P}{\partial y^2}=-2"

"D(x, y)=\begin{vmatrix}
 -2 & -1 \\
 -1 & -2
\end{vmatrix}=4-1=3"

"\dfrac{\partial^2 P}{\partial x^2}(100,200)=-2<0"

"D(100, 200)=3>0"

It can be stated that "(100,200)" is a relative maximum.


"P(100,200)=400(100)+500(200)"

"-(100)^2-(200)^2-100(200)-20000=50000"

The Profit has the maximum with value of "50000" at "x=100" units and "y=200" units.


The Profit has no minimum.